# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v1

@@ 2 Preliminaries
3 Diffusion Latent Beam Search
We first provide a unified view of existing inference-time alignment methods through several practical approximations of optimal drift $\mathbf{u}^{*}(\boldsymbol{\nu}_{t},t)$ (Section 3.1). To mitigate errors from those approximations, we propose a novel inference-time search algorithm, diffusion latent beam search with lookahead estimator (Section 3.2).
3.1 A Unified View for Practical Approximations
While Equation 8 has a relatively tractable form, it is still computationally expensive, since the expectation requires full diffusion sampling to evaluate latent at each time step, as well as facing numerical instability. Previous alignment methods rely on multiple-step practical approximations.
Step. 1: Jensen’s Inequality   First, when assuming $\frac{r^{\prime}(\boldsymbol{\nu}_{T})}{\lambda}$ is almost deterministic (this might hold when $t\rightarrow T$), Jensen’s inequality yields the following approximation by exchanging $\log$ and $\mathbb{E}[\cdot]$, which can be considered as a certain form of classifier guidance (Dhariwal & Nichol, 2021):
Step. 2: Tweedie’s Formula   To avoid the computationally expensive expectation, the expected reward is further approximated as $\mathbb{E}_{p^{\text{pre}}}[r^{\prime}(\boldsymbol{\nu}_{T})|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]\approx r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ where $\hat{\boldsymbol{\nu}}_{T|t}\approx\mathbb{E}_{p^{\text{pre}}}[\boldsymbol{\nu}_{T}|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]$ is a one-step approximation of posterior mean (Chung et al., 2023), which can be calculated only with the current latent $\boldsymbol{\nu}_{t}$ without full diffusion path. Therefore, the optimal drift term can be seen as solely depending on the current time step $t$:
Such a computationally tractable drift term has been leveraged for previous inference-time alignment methods via approximate guidance or twisted sequential Monte Carlo (SMC) (Wu et al., 2024b). However, as the approximated posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$in intermediate steps is noisy, evaluation with the reward function for clean data $r^{\prime}(\cdot)$ may not provide a reliable signal (Liang et al., 2024). Moreover, Equation 10 requires the reward gradient, which is not applicable to non-differentiable reward, such as AI feedback, and is also not suitable for the modality whose reward gradient imposes a huge computational cost in practice, such as video.
Step. 3: Converting Reward Gradient into argmax   The usage of reward gradient can be converted into $\operatorname{argmax}$ operator (Huang et al., 2024a; Li et al., 2024b; Bansal et al., 2023). The intuition here is since the optimal drift in Equation 10 induces the diffusion latent to the direction where it maximizes the reward, we replace such a maximization with a zeroth-order search. The SDE is approximated as:
$d\boldsymbol{\nu}=\bar{\textbf{f}}(\boldsymbol{\nu},t)dt+g(t)d\textbf{w}^{*}~{}~{}\text{where}~{}~{}d\textbf{w}^{*}=\underset{d\textbf{w}}{\operatorname{argmax}}~{}r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t}).$ (11)
Note that the current diffusion latents $\boldsymbol{\nu}_{t}$ and posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$ are sampled by following the standard Wiener process $d\textbf{w}$. This approximation is leveraged for inference-time alignment via greedy search (Huang et al., 2024a; Li et al., 2024b) or SMC (Singhal et al., 2025) of diffusion latents. However, greedy search can result in sub-optimal generation affected by inaccurate reward estimate $r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ due to its noisy input. Moreover, for the high-dimensional domain, such as video generation, it can be challenging to get an accurate density ratio term required in SMC.
Figure 3: 2D-histogram and correlation between reward functions for perceptual video quality (Huang et al., 2024b) and AI feedback from Gemini (Gemini Team, 2023). A single reward (e.g., subject consistency; blue) is often not aligned well with a preference from Gemini, which happens for all the prompt sets with different degree of dynamics. The calibrated reward, a linear combination of perceptual metrics via brute-force search (green), achieves the best Pearson correlation coefficient in all settings (statistically significant with $p<0.01$).
3.2 Mitigating Approximation Errors via Beam Search
Existing practical algorithms based on these three approximations, such as greedy search (Huang et al., 2024a; Li et al., 2024b), fall into sub-optimal generation due to the erroneous reward evaluation with a noisy estimate of the posterior mean (Chung et al., 2023), and $\operatorname{argmax}$ operator based on them. To resolve the error accumulation, we propose a simple yet robust modification, diffusion latent beam search (DLBS) with lookahead estimator. To clearly describe the practical implementation, we use the notation of discrete-time diffusion process in the rest of the section.
Practical Implementation   We summarize the detailed sampling procedure of DLBS in Algorithm 1. For the diffusion sampler, we use stochastic DDIM (Song et al., 2021a) with a decreasing sequence $\{\alpha_{t}\}_{t=1}^{T}\in(0,1]^{T}$, noise level $\eta$, and noise schedule $\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t-1}/\alpha_{t}}$, which is equivalent to DDPM (Ho et al., 2020) when $\eta=1.0$. We initialize $B$ latent beams from the Gaussian distribution (Line 1.1), sample $K$ latents per beam in next time step (Line 1.11), and then compute the one-step estimation of the posterior mean (Line 1.13). DLBS evaluates the estimator of posterior mean $\hat{\mathbf{z}}_{0|t-1}$ with reward function (Line 1.18) and selects Top-$B$-rewarded latent beams instead of Top-$1$ (i.e., $\operatorname{argmax}$) from $KB$ candidates (Line 1.19), which is iterated over entire reverse process from $t=T$ to $t=0$. DLBS can possess latent beams more widely than greedy search under the same budget, which mitigates error propagation due to the approximated diffusion latent evaluation.
@@ 3.2 Mitigating Approximation Errors via Beam Search
Lookahead Estimator   The other source of approximation errors than $\operatorname{argmax}$ operator is a one-step estimator of the posterior mean $\hat{\mathbf{z}}_{0|t-1}$ from Tweedie’s formula, which is still noisy, especially in earlier time steps and leads to inaccurate reward evaluation. To reduce errors in reward evaluation, we propose a lookahead (LA) estimator $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$, which is estimated by running $T^{\prime}$-step deterministic DDIM $(1<T^{\prime}<<T)$ while equally interpolating the rest of time steps from the current latent $\mathbf{z}_{t-1}$ to $\mathbf{z}_{0}$ (Algorithm 2). While requiring additional denoising steps, its cost is almost the same as naive DLBS because most computational costs come from when we decode $\textbf{z}_{0}$ (i.e. reward evaluation). This more “accurate” estimator significantly improves the performance even with $T^{\prime}=2$ or $3$ (Section 5.2).
@@ 4 Calibrating Reward to Preference Feedback
One of the most useful assessments for generative models is human evaluation, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce the time and cost is to leverage AI feedback from VLMs (Wu et al., 2024d), which has been shown to align with human judgment on video quality to some extent (Na et al., 2024; Furuta et al., 2024b). In this work, we assume that the VLM evaluation works as an oracle and we align model outputs to the preference of VLMs, which is reasonable due to their capability and the cost to be saved. Our qualitative evaluation also confirms that the high-rating video by VLMs is good for us. It is possible to replace VLMs with humans in our framework if the cost allows.
However, because alignment via inference-time search requires massive reward evaluation queries, we still need to build more tractable proxy rewards not relying on humans or external VLM APIs. The question here is what kind of metrics for perceptual video quality can improve the feedback from VLMs. Because the criteria of videos preferable to humans are multi-objective, maximizing a single metric may lead to undesirable generation due to over-optimization. In this section, we first review the possible video quality metrics (Section 4.1), evaluate the Pearson correlation between these and VLM feedback score, and then propose a reward calibration (Section 4.2), aiming to align the existing video rewards to VLMs by considering their weighted linear combination through the brute-force search of coefficients.
@@
Figure 5: Comparison among diffusion latent beam search (DLBS), best-of-N (BoN) and greedy search (GS). We measure the performance in terms of a combinational reward calibrated to Gemini (above) and GPT-4o (below). DLBS improves all the calibrated reward the best as the search budget $KB$ increases (especially $KB=16,32$) while BoN and GS in some cases eventually slows down or saturates the performance. Notably, LA estimator ($KB=8,T^{\prime}=6$) is comparable to or even outperforming DLBS ($KB=32$).
4.1 Metric Reward for Perceptual Video Quality
Following Huang et al. (2024b), we select six base reward functions for perceptual video quality (see Appendix B):
Subject Consistency quantifies how consistently the subject appears across video frames with DINO (Caron et al., 2021).
Motion Smoothness leverages the motion prior in AMT (Li et al., 2023) to evaluate whether the generated video’s motion is smooth and physically plausible.
Dynamic Degree quantifies the overall magnitude of dynamic object movement by estimating optical flow (Teed & Deng, 2020) for each pair of consecutive frames.
Aesthetic Quality measures compositional rules, color harmony, and the overall artistic merit of each video frame with LAION aesthetic predictor (LAION-AI, 2022).
Imaging Quality assesses low-level distortions (e.g., over-exposure, noise, blur) in each frame with MUSIQ predictor (Ke et al., 2021).
Text-Video Consistency captures how closely a content in a video aligns with a prompt with ViCLIP (Xu et al., 2021).
Reward Calibration   To reflect the multi-dimensional aspect of preferable videos, we model the calibrated reward function $r^{*}(\cdot,\cdot)$ as a weighted linear combination of video quality metrics: $r^{*}(\mathbf{x}_{0},\mathbf{c}):=\sum_{i=1}^{M}w_{i}r_{i}(\mathbf{x}_{0},\mathbf{c})$. The coefficient $w_{i}$ is determined by maximizing the Pearson correlation with preference feedback. We heuristically conduct a brute-force search within a reasonable range (Section 4.2).
Experimental Setup   We leverage GPT-4o (OpenAI, 2023) and Gemini-1.5 (Gemini Team, 2023) as automated raters for generated videos. We provide a text prompt and generated video as inputs, instructing VLMs to assign discrete scores (1–10) based on overall visual quality (including clarity, resolution, brightness, and aesthetic appeal), the appropriateness of motion for either static or dynamic scenes, the smoothness and consistency of shapes and motions, and the degree of alignment with a prompt (see Appendix E).
We select four prompt sets from two distinct dataset (see Appendix C). DEVIL (Liao et al., 2024) classifies its prompts into five categories depending on the degree of movement, each further divided by subject type (e.g., cat, horse, truck, nature, etc.). We focus on three of five categories (static, medium, and high) and select one prompt at random from each subject-subdivision within a chosen category. We also draw 30 random captions from the test split of MSRVTT (Xu et al., 2016), widely used as a video benchmark.
To examine the correlation among AI feedback and perceptual quality metrics, we generate 64 videos per prompt from pre-trained Latte (Ma et al., 2024) using DDIM sampler with $T=50$ and $\eta=0.0$. We also prepare candidates of the calibrated reward by choosing the combination of weights $w_{i}\in\{0.0,0.25,0.5,0.75,1.0\}$ and use those later to rank them based on the correlation with AI feedback.
@@ 4.2 Correlation and Reward Calibration
Figure 3 illustrates the 2D-histogram and corresponding correlation between each individual metric and feedback from Gemini (see Appendix F for GPT-4o). Relying on a single metric often yields low correlation, which supports the multifaceted nature of perceptual video quality. In addition, the relative importance of each metric depends on the movement in the prompts; in highly-dynamic DEVIL-high, dynamic degree correlates more strongly with VLMs than consistency metrics. Conversely, in less-dynamic DEVIL-medium or DEVIL-static, subject consistency and motion smoothness play more prominent roles. Because the aesthetic score focuses on frame-by-frame visual quality, it tends to exhibit a stronger correlation with VLM in low-motion scenarios, while rapid movements and frequent transitions often introduce motion blur or abrupt changes in composition, reducing the frame-level aesthetic quality and thus weakening its correlation with VLMs.
Notably, a weighted linear combination of these metrics yields the highest correlation with Gemini (Figure 3, green). Among brute-force candidates, we select the best coefficients in terms of correlation with Gemini per a set of prompts in Figure 4, which differ based on the degree of movement in the prompts. The dynamic degree and motion smoothness exhibit pronounced variation among DEVIL-high, -medium, and -static, and the weights of other factors, such as text-video consistency, subject consistency, and aesthetic considerations, are also adjusted to the prompts accordingly. This underscores the importance of appropriately weighting multiple evaluation criteria through the reward calibration.
@@ 5 Inference-Time Text-to-Video Alignment
Experimental Setup  We compare following inference-time search methods with a noise level $\eta=1.0$ for DDIM:
@@
  • DLBS:  Given the budget $KB$, we sweep possible combinations in terms of multiple of 2 (e.g., $K=4,B=4$ or $K=8,B=2$), and then report the best results except for the case with $K=1$ and $B=1$.
@@
  • DLBS-LA:  We combine DLBS with lookahead estimator from 6-step deterministic DDIM. Due to the resources, we have experiments with $KB=8$.
We use the same prompt sets as in Section 4, and Gemini-/GPT-calibrated rewards in Section 4.2. Note that the range of values may be different from each other as they have different coefficients. We test the performance when scaling (1) search budget $KB$ (Section 5.1), (2) LA steps $T^{\prime}$ for reward estimate (Section 5.2), and (3) denoising steps (Section 5.3). We further analyze whether DLBS maximizing reward calibrated to VLMs can actually improve VLM evaluation (Section 5.4), and how the diversity of generated samples is in the inference-time search (Section 5.5).
5.1 Scaling Search Budget
@@
Figure 5 measures the combinational reward calibrated to Gemini (above) and GPT-4o (below) while increasing the search budget $KB\in\{1,2,4,8,16,32\}$. DLBS improves all the calibrated rewards the best as $KB$ increases (especially $KB=16,32$) while BoN and GS in some cases eventually slow down or saturate the performance. See Appendix H for the results with $KB=64$ in DEVIL-medium where we still observe the improvement.
@@
Figure 6 demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beam $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we have a peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS help search for the best outputs robustly.
Figure 7: Comparison among different LA steps $T^{\prime}$ on MSRVTT-test ($KB=8$). The performance improves, as increasing LA steps. Even $T^{\prime}=2$ significantly outperforms the posterior mean.
5.2 Scaling Lookahead Steps for Reward Estimate
We scale up LA steps $T^{\prime}$ to obtain an accurate reward estimate. We use MSRVTT-test ($KB=8$) and Gemini reward for experiments. Figure 7 shows that as increasing LA steps, the performance improves more. Even $T^{\prime}=2,3$ significantly outperforms the posterior mean, which is often used in prior works (Kim et al., 2024; Huang et al., 2024a; Singhal et al., 2025). This is because the sub-optimal performance of inference-time search comes from the approximation errors and LA estimator can notably reduce them. Moreover, as shown in Figure 5 (above), DLBS-LA $(KB=8,T^{\prime}=6)$ achieves comparable or even outperforming results with DLBS $(KB=32)$. It is quite beneficial to spend a computation to estimate reward accurately.
5.3 Scaling Denoising Steps
Figure 8: (Left) Scaling the denoising steps $T$. DLBS benefits from larger $T$ more than BoN ($K=1$) or GS ($B=1$). (Right) Range of denoising steps $t\in[50,0]$ to apply search methods. While Kim et al. (2024) apply GS in the first 5–10 steps, DLBS over the entire diffusion steps yields the largest improvement.
The last axis is denoising steps; in general, the more denoising steps are, the finer the video generated is (Ho et al., 2020). We investigate (1) whether this also holds in DLBS, (2) how the improvement is compared to other axes, and (3) when we should run DLBS, while Kim et al. (2024) only apply GS in the first 5–10 steps.
Figure 8 (left) shows the performance when increasing the number of denoising steps $T$. Since DDIM exhibits fast convergence (Song et al., 2021a), BoN with larger $T$ does not improve the reward much. DLBS improves the performance when scaling denoising steps to $T=200$ more than BoN or GS, which implies that DLBS benefits from larger computation as well as the other axes. As shown in Figure 5 (MSRVTT-test), the reward is comparable to the one in $KB=16$, although the range of improvement by denoising steps is smaller than the one from the LA estimator. The last point is how many times we should iterate DLBS. As opposed to Kim et al. (2024), Figure 8 (right) shows that applying DLBS for all the steps is much better than that for limited steps. Scaling the timing when we employ DLBS realizes better alignment on a calibrated reward, and the number of time steps may also improve the performance, while the LA estimator is a priority.
5.4 Evaluation with AI Feedback

### v2

@@ 3.2 Mitigating Approximation Errors via Beam Search
Lookahead Estimator   The other source of approximation errors than the $\operatorname{argmax}$ operator is a one-step estimator of the posterior mean $\hat{\mathbf{z}}_{0|t-1}$ from Tweedie’s formula, which is still noisy, especially in earlier time steps, and leads to inaccurate reward evaluation. To reduce errors in reward evaluation, we propose a lookahead (LA) estimator $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$, which is estimated by running $T^{\prime}$-step deterministic DDIM $(1<T^{\prime}<<T)$ while equally interpolating the rest of time steps from the current latent $\mathbf{z}_{t-1}$ to $\mathbf{z}_{0}$ (Algorithm 2). While requiring additional denoising steps, its cost is almost the same as naive DLBS because most computational costs come from when we decode $\textbf{z}_{0}$ (i.e., reward evaluation). Theoretically, enlarging the lookahead steps $T^{\prime}$ monotonically tightens the upper bound on the reward-approximation error (see Appendix E). Empirically, a modest horizon ($T^{\prime}=2,3,6$) delivers substantial search improvements (Figure 8; Left), but the marginal gains saturate, so pushing $T^{\prime}$ further yields little additional benefit (see Appendix L.4).
@@ 4 Calibrating Reward to Preference Feedback
Human evaluation is one of the most valuable assessments for generative models, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce the time and cost is to leverage AI feedback from VLMs [43], which has been shown to modestly align with human judgment on video quality [44, 45] (see Appendix J). In this work, we assume that the VLM evaluation works as an oracle, and we align model outputs with the preference of VLMs, which is reasonable due to their capability and the cost to be saved. Our qualitative and quantitative evaluations also confirm that the high-rated video by VLMs is generally good for us.
However, because alignment via inference-time search requires massive reward evaluation queries, we still need to build more tractable proxy rewards that do not rely on humans or external VLM APIs. The question here is what metrics for perceptual video quality can improve the feedback from VLMs. Because the criteria of videos preferred by humans are multi-objective, maximizing a single metric may lead to undesirable generation due to over-optimization. For instance, focusing exclusively on temporal consistency or frame-by-frame quality metrics can unintentionally reduce the video’s motion magnitude. In this section, we first review the possible video quality metrics (Section 4.1), evaluate the Pearson correlation between these and VLM feedback score, and then propose a reward calibration (Section 4.2), aiming to align the existing video rewards to VLMs by considering their weighted linear combination through the brute-force search of coefficients.
Figure 3: 2D-histogram and correlation between reward functions for perceptual video quality [12] and AI feedback from Gemini [14]. A single reward (e.g., subject consistency; blue) is often not aligned well with a preference from Gemini, which happens for all the prompt sets with different degrees of dynamics. The calibrated reward, a linear combination of perceptual metrics via brute-force search (green), achieves the best Pearson correlation coefficient in all settings (statistically significant with $p<0.01$).
4.1 Metric Reward for Perceptual Video Quality
Following Huang et al. [12], we select six base reward functions for perceptual video quality (see Appendix C):
  • Subject Consistency quantifies how consistently the subject appears across video frames with DINO [46].
  • Motion Smoothness leverages the motion prior in AMT [47] to evaluate whether the generated video’s motion is smooth and physically plausible.
  • Dynamic Degree quantifies the overall magnitude of dynamic object movement by estimating optical flow [48] for each pair of consecutive frames.
  • Aesthetic Quality measures compositional rules, color harmony, and the overall artistic merit of each video frame with LAION aesthetic predictor [49].
  • Imaging Quality assesses low-level distortions (e.g., over-exposure, noise, blur) in each frame with MUSIQ predictor [50].
  • Text-Video Consistency captures how closely the content in a video aligns with a prompt with ViCLIP [51].
@@ 4.1 Metric Reward for Perceptual Video Quality
Reward Calibration  To reflect the multi-dimensional aspect of preferred videos, we model the calibrated reward function $r^{*}(\cdot,\cdot)$ as a weighted linear combination of video quality metrics: $r^{*}(\mathbf{x}_{0},\mathbf{c}):=\sum_{i=1}^{M}w_{i}r_{i}(\mathbf{x}_{0},\mathbf{c})/\sum_{i=1}^{M}w_{i}$. The coefficient $w_{i}$ is determined by maximizing the Pearson correlation with preference feedback. We heuristically conduct a brute-force search within a reasonable range (Section 4.2).
Experimental Setup  We leverage Gemini-1.5 [14] and GPT-4o [13] as automated raters for generated videos. We provide a prompt and generated video as inputs, instructing VLMs to assign discrete scores (from 1 to 10) based on overall visual quality (e.g., clarity, resolution, brightness, and aesthetic appeal), the appropriateness of motion for either static or dynamic scenes, the smoothness and consistency of shapes and motions, and the degree of alignment with a prompt (see Appendix H).
We select four prompt sets from two distinct datasets (see Appendix F). DEVIL [52] classifies its prompts into five categories depending on the degree of movement, each further divided by subject type (e.g., cat, horse, truck, nature, etc.). We focus on three of five categories (high, medium, and static) and select one prompt randomly from each subject-subdivision within a chosen category. We also draw 30 random captions from the test split of MSRVTT [53], widely used as a video benchmark.
We generate 64 videos per prompt from pre-trained Latte [18] using the DDIM sampler with $T=50$ and $\eta=0.0$ to examine the correlation among AI feedback and perceptual quality metrics. We also prepare candidates for the calibrated reward by choosing the combination of weights $w_{i}\in\{0.0,0.25,0.5,0.75,1.0\}$ and use those later to rank them based on the correlation with AI feedback.
@@ 4.2 Correlation and Reward Calibration
Figure 3 illustrates the 2D-histogram and corresponding correlation between each metric and feedback from Gemini. Relying on a single metric often yields low correlation, which supports the multifaceted nature of perceptual video quality. See Appendix I.1 for further results, where we can see that the relative importance of each metric depends on the movement in the prompts; in highly dynamic DEVIL-high, the dynamic degree correlates more strongly with VLMs than consistency metrics. Conversely, subject consistency and motion smoothness play more prominent roles in less-dynamic DEVIL-medium or DEVIL-static. Because the aesthetic score focuses on frame-by-frame visual quality, it tends to correlate strongly with VLM in low-motion scenarios. Rapid movements and frequent transitions often introduce motion blur or abrupt changes in composition, reducing the frame-level aesthetic quality and thus weakening its correlation with VLMs.
Reward calibration, a weighted linear combination of these metrics, yields the highest correlation with Gemini (Figure 3, green). We select the best coefficients among brute-force candidates, based on the correlation with Gemini, for each set of prompts that have a different description of the degree of movement (Figure 4). The dynamic degree and motion smoothness exhibit pronounced variation among DEVIL-high, -medium, and -static, and the weights of other factors, such as text-video consistency, subject consistency, and aesthetic considerations, are also adjusted accordingly. In addition, Appendix K presents results from best-of-64 sampling with a single metric or calibrated reward, where a single metric often leads to over-optimization. This highlights the importance of reward calibration, appropriately weighting multiple criteria, as aligning with a request in the prompt.
Figure 5: (Left) Comparison among diffusion latent beam search (DLBS), best-of-N (BoN), and greedy search (GS). We measure the performance in terms of a combinational reward calibrated to Gemini. DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$). (Right) Optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget.
@@ 5 Inference-Time Text-to-Video Alignment
Experimental Setup  We use the same prompts and Gemini-/GPT-calibrated rewards as in Section 4. We compare the following inference-time search methods with a noise level $\eta=1.0$ for DDIM:
@@
  • DLBS:  Given the budget $KB$, we sweep possible combinations in terms of power of 2 (e.g., $K=8,B=2$), and report the best results except for the case with $K=1$ and $B=1$.
@@
  • DLBS-LA:  We combine DLBS with a lookahead estimator from the 6-step deterministic DDIM.
We normalize the rewards with min/max values so that the range is around $[0,1]$. Our experiments aim to assess: (1) scaling the search budget and computational costs for efficient resource allocation (Section 5.1); (2) evaluating alignment performance with feedback from humans and VLMs (Section 5.2); (3) assessing scalability to capable SoTA models (Section 5.3); (4) performing detailed ablations on LA steps $T^{\prime}$, and robustness to diverse and complex prompts (Section 5.4); (5) quantitative analysis on the diversity of generated video (Appendix L.7).
5.1 Scaling Search Budget and Computational Cost
@@
Figure 5 (Left) measures the combinatorial reward calibrated to Gemini while increasing the search budget $KB\in\{1,2,4,8,16,32\}$. DLBS improves all the calibrated rewards the best as $KB$ increases (especially $KB=16,32$), while BoN and GS eventually slow down or saturate the performance in some cases. See Appendix L.1 for results with GPT calibrated reward and Appendix L.2 for the results with $KB=64$, where we still observe the improvement.
@@
Figure 5 (Right) demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly. See Appendix L.3 for further results.
We also analyze how alignment performance scales with different DLBS configurations under fixed computational budgets, using function evaluations (NFE) and wall-clock time as cost measures. As shown in Figure 6 (Left), DLBS consistently outperforms BoN and GS across all budgets, demonstrating superior efficiency in utilizing compute. Adding the LA estimator further amplifies this advantage, offering substantial performance gains with minimal overhead. In contrast, increasing the number of diffusion steps $T$ (DLBS-T; $KB=8$, $T\in\{100,200\}$) results in only marginal improvements despite the higher computational cost. Our results suggest a clear strategy for inference-time budget allocation: prioritize enabling the LA estimator and increasing the search budget $KB$ leads to substantial performance gains, while increasing the diffusion steps $T$ provides limited benefit relative to its computational cost.
Figure 6: (Left) DLBS achieves alignment performance gains more efficiently than BoN and GS under the same number of function evaluations (NFE) or execution time. Increasing the search budget provides larger improvements under an equivalent computational cost than scaling the number of diffusion steps (DLBS-T; $KB=8,T=100,200$). Employing the LA estimator (DLBS-LA) further amplifies these gains, only with marginal overhead, yielding remarkably better efficiency than BoN or GS. (Right) DLBS and DLBS-LA help the latest SoTA models, CogVideoX-5B [19] and Wan 2.1-14B [20], improve the generated video quality.
5.2 Evaluation with AI and Human Feedback